Teacher gave us the answers which i included, but i need to know how to work the problems for the test! I promise to rate the answer!Paired Samples Statistics Std. Error Mean Results for husbands satisfaction Mean Std. Deviation 4.33 4.83 Pair h sat t 31.78 30.28 76 76 50 h sat t2 t1- time 1 t2time 2 Paired Samples Correlations Si Correlation 68 Pair 1 h sat t1 & h sat t2 76 Paired Samples Test Paired Diffe rences 95% Confidence Interval of the Difference Std. Error Mean Std. Deviation Mean Lower df Sig. (2-tailed) Pair 1 hsat t1 h satt2 1.50 3.68 42 75 84. Compute the standard error of the mean for husbands at time 2. 85. Compute the pooled variance 86. Compute the lower end of the 95% confidence interval 87. Compute the upper end of the 95% confidence interval ÍgyJ 0.55% ss) 긔.033
Answer #1

84) Standard error of mean for husband time 2 is s/V(n) 7l = 4.83/\sqrt(76) = 4.83/8.72= 0.554

85) Pooled variance

The following information has been provided:

s1​=4.33, n1​=76, s2​=4.83 and n2​=76

The following formula is used to compute the pooled variance:

sp^2​=(n1​−1)s1^2​+(n2​−1)s2^2/(​​n1​+n2​−2)

Plugging in the corresponding values in the formula above, we get that sp^2​=(76−1)(4.33)^2+(76−1)(4.83)^2/(76+76−2)​=21.0389

Now that we have the pooled variance, we can compute the pooled standard deviation by simply taking the squared root of the value obtained for the variance.

The following pooled standard deviation is obtained:

sp​=sqrt(21.0389) ​=4.5868

86) degrees of freedom= n-1= 76-1=75

alpha=0.05

t critical value= 1.99

95% confidence interval

lower end= mean-tc*Standard error= 1.5-1.99*0.42=0.66

87) Upper end= 1.5+1.99*0.42= 2.34